http://pleiad.dcc.uchile.cl/

CV vramiro

Victor Ramiro August 19, 2008

http://pleiad.dcc.uchile.cl/ http://pleiad.dcc.uchile.cl/

Contents

0.1 About Me . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 0.2 Short Research Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 0.3 Publications . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 0.4 Teaching Assistant . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 0.5 Other Stuff . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

3/6

http://pleiad.dcc.uchile.cl/

4/6

http://pleiad.dcc.uchile.cl/

About Me  CONTENTS

img /home/v/pleiad/www/data/media/people/vramiro.png Victor Ramiro

Master Student vramiro@NOSPAM-dcc.uchile.cl PLEIAD Lab Computer Science Department (DCC) Blanco Encalada 2120 University of Chile Santiago, Chile

0.1 About Me

My name is Victor Ramiro (yeah, Ramiro is my surname!), I’m studying at the Computer Science Department of the University of Chile. Currently I’m working on my Master Thesis promoted by Prof. Eric Tanter.

0.2 Short Research Description

We study a conceptual framework to plug proximity as a first-class concept in the AmbientTalk language. This means developers will be able to create and share different notions of proximity and evaluate them with the context information at the service discovery stage. The contribution of this work is to support a wide variety of proximity notions such that the device interactions can be scaled and tailored to the content of interest level.

0.3 Publications

The following documents are included by the contributing authors as a means to ensure timely dissemination of scholarly and technical work on a non-commercial basis. Copyright and all rights therein are maintained by the authors or by other copyright holders, notwithstanding that they have offered their works here electronically. It is understood that all persons copying this information will adhere to the terms and constraints invoked by each author’s copyright. These works may not be reposted without the explicit permission of the copyright holder.

1 <html>

2

3 <s c r i p t>l o a d b i b t e x ( ’ vramiro ’ )</s c r i p t>
4 <div id=”b i b t e x ”>Loading b i b t e x i n f o …</div>

5

6 </html>

0.4 Teaching Assistant

I’ve been Teaching Assistant in the following Courses:

here

• Fall 2005 - Prof. Alejandro Hevia

• Fall 2006 - Prof. Valeria Herskovic

• Fall 2007 - Prof. Juan Alvarez

• Fall 2004 - Prof. Julio Salas, Hector Mora, Cristian Cespedes

• Spring 2004 - Prof. Julio Salas, Hector Mora

• Summer 2004 - Prof. Julio Salas

• Spring 2005 - Prof. Hector Mora

here

• Spring 2005 - Prof. Sergio Ochoa

5/6

http://pleiad.dcc.uchile.cl/

Other Stuff  CONTENTS

• Spring 2006 - Prof. Cecilia Bastarrica here

• Fall 2006 - Prof. Luis mateu

• Fall 2007 - Prof. Eric Tanter here

• Fall 2007 - Prof. José Miguel Piquer

0.5 Other Stuff

• The AmbientTalk Group

• Personal Webpage

• My Linkedin Network

6/6